A spring of natural lengthb19inches stretches 1.5 inches under a weight of pounds .find the work done in stretching the spring
1
Expert's answer
2020-06-15T19:25:10-0400

Let, the equilibrium point is at origin i.e "x_i=0"

Information provided,

Natural length of spring "l_0=19" inches.

Changes in the length of spring from natural length is "x=1.5\: inch=0.0381m"

Force acting on the spring is 1 pound weight "= 4.45N."

As, magnitude of force is

"F=kx\\\implies
4.5=k(0.0381)\implies k=116.67N/m"

Since, work done in spring is the potential energy stored in the spring due to stretching or compressing it.

Since,

"W=\Delta PE=\int_{x_i}^{x}kxdx=\frac{1}{2}kx^2-\frac{1}{2}kx_i^2"

But in this case, "x_i=0"

Thus,

"W=\frac{1}{2}kx^2"

On plugin the value we get,

"W=8.5\times 10^{-2}J"